A share trader has a policy of selling a share if its share price drops more than one-third of its

purchase price. In January, the investor bought shares in Netron (an electronic company) for 90c

per share. The table below shows the price of the Netron share at the end of each month.

Month Jan Feb Mar Apr May Jun Jul

Share price 90 82 78 80 74 76 70

The three-period moving average of the share price in April is

1. 83:3

2. 80

3. 75:6

4. 76:7

5. 77:33


1
Expert's answer
2021-08-18T07:43:32-0400

Three-period moving average "= \frac{Total}{count}"

"= \frac{78+ 80+ 74 }{3}\\
= \frac{232 }{3}\\
=77.33"